# A highly efficient single-loop smoothing damped Newton method for large-scale bilevel hyperparameter optimization of SVC

### v1

@@ 7 Conclusion
In this paper, we develop an efficient algorithm to solve the (MPEC) derived from the bilevel optimization model for hyperparameter selection for SVC. Specifically, we smoothed the complementarity constraints in the (MPEC). We analyzed theoretical properties of the resulting smoothing problem (NLPEC), including its feasible sets, Jacobian matrix and second order sufficient condition. Then we applied the damped Newton method studied in [31] to solve (NLPEC), in which BICG-STAB algorithm was used to solve KKT equation system. Under certain assumption, the proposed smoothing damped Newton method (SDNM) enjoys a quadratic convergence rate. Finally, extensive numerical results on the datasets from the LIBSVM library verified the efficiency of SDNM over almost all the datasets used in this paper comparing with the other two methods. SDNM demonstrates high efficiency in practice with superior generalization performance on classification prediction and solving the over-fitting problem, and is about twice as fast as the better one of the existing approaches (fmincon and SGRM). For an example with sample data 450 and feature number 300, it can return a solution of hyperparameter in 10 seconds, which is over 20 times faster than SGRM and over 3 times faster than fmincon. SDNM applied the smoothing function and the damped Newton method to the nonsmooth bilevel problem with $l_{1}$ loss and $l_{0/1}$ loss, which has the potential to deal with other nonsmooth hyperparameter classification problems in SVM. These topics will be investigated further in the near future.
Appendix A
Table 5: Computational results of SDNM on different smoothing parameters
No. Dataset $\epsilon$ $C$ $t$ $E_{tr}$ $E_{CV}$ $A_{2}$ $(k,iter)$
1 diabetes $10^{-5}$ 3.86 5.6 $\mathbf{23.3}$ 23.3 991.19 $(10,741)$
$10^{-4}$ 3.65 7.3 $\mathbf{23.3}$ 23.3 990.78 $(12,812)$
$10^{-3}$ 3.08 3.7 24.1 $\mathbf{23.0}$ 720.06 $(7,441)$
$10^{-2}$ 3.21 15.5 24.6 $\mathbf{23.0}$ 904.31 $(26,2050)$
2 heart $10^{-5}$ 0.45 0.9 $\mathbf{15.8}$ $\mathbf{18.0}$ 5.51 $(6,1013)$
$10^{-4}$ 0.42 1.1 $\mathbf{15.8}$ $\mathbf{18.0}$ 18.39 $(7,1058)$
$10^{-3}$ 0.60 0.9 16.7 $\mathbf{18.0}$ 10.11 $(6,897)$
$10^{-2}$ 0.72 0.9 $\mathbf{15.8}$ 18.7 1.57 $(7,950)$
3 australian $10^{-5}$ 2.72 4.3 14.6 13.7 21.25 $(5,682)$
$10^{-4}$ 0.80 10.4 $\mathbf{13.3}$ 14.3 192.43 $(11,1780)$
$10^{-3}$ 3.62 3.2 14.9 14.0 9.98 $(4,472)$
$10^{-2}$ 1.68 6.1 14.9 $\mathbf{13.3}$ 41.19 $(7,1029)$
4 german.number $10^{-5}$ 15.00 1.5 $\mathbf{25.3}$ $\mathbf{26.7}$ 965.07 $(5,52)$
$10^{-4}$ 15.01 6.2 $\mathbf{25.3}$ $\mathbf{26.7}$ 867.54 $(14,576)$
$10^{-3}$ 15.00 1.6 $\mathbf{25.3}$ $\mathbf{26.7}$ 950.78 $(5,57)$
$10^{-2}$ 15.05 6.9 $\mathbf{25.3}$ $\mathbf{26.7}$ 877.57 $(13,624)$
5 ionosphere $10^{-5}$ 8.45 2.3 $\mathbf{2.7}$ $\mathbf{26.7}$ 747.04 $(6,366)$
$10^{-4}$ 6.37 2.6 $\mathbf{2.7}$ $\mathbf{26.7}$ 733.99 $(7,466)$
$10^{-3}$ 8.26 2.4 $\mathbf{2.7}$ $\mathbf{26.7}$ 640.83 $(6,433)$
$10^{-2}$ 7.39 3.5 $\mathbf{2.7}$ $\mathbf{26.7}$ 885.00 $(8,659)$
6 sonar $10^{-5}$ 3.72 0.6 $\mathbf{27.6}$ $\mathbf{28.7}$ 225.52 $(5,409)$
$10^{-4}$ 4.61 0.7 $\mathbf{27.6}$ 30.7 411.16 $(8,360)$
$10^{-3}$ 6.46 0.9 $\mathbf{27.6}$ 30.7 447.09 $(9,353)$
$10^{-2}$ 4.83 1.5 $\mathbf{27.6}$ $\mathbf{28.7}$ 447.29 $(20,363)$
7 phishing $10^{-5}$ 3.56 5.7 $\mathbf{6.2}$ $\mathbf{7.0}$ 2339.57 $(10,724)$
$10^{-4}$ 3.56 5.3 $\mathbf{6.2}$ $\mathbf{7.0}$ 2223.30 $(11,631)$
$10^{-3}$ 0.56 4.2 7.1 11.0 11.95 $(9,493)$
$10^{-2}$ 0.74 10.6 7.1 11.7 357.77 $(18,1416)$
8 w1a $10^{-5}$ 0.01 9.6 2.8 $\mathbf{2.7}$ 35.35 $(9,472)$
$10^{-4}$ 0.05 9.7 $\mathbf{2.5}$ 3.1 39.18 $(8,541)$
$10^{-3}$ 0.03 8.9 3.1 3.8 43.87 $(8,458)$
$10^{-2}$ 0.08 28.2 $\mathbf{2.5}$ $\mathbf{2.7}$ 47.77 $(17,1916)$
9 w2a $10^{-5}$ 3.69 8.6 $\mathbf{3.2}$ 3.6 1014.79 $(5,277)$
$10^{-4}$ 2.84 8.3 $\mathbf{3.2}$ $\mathbf{3.3}$ 1141.97 $(5,350)$
$10^{-3}$ 3.49 8.8 $\mathbf{3.2}$ 3.6 987.18 $(5,323)$
$10^{-2}$ 3.30 6.7 $\mathbf{3.2}$ 3.8 865.45 $(4,232)$
10 w3a $10^{-5}$ 0.29 11.3 2.1 2.9 149.90 $(7,730)$
$10^{-4}$ 0.29 16.2 $\mathbf{2.0}$ 2.9 715.60 $(9,1131)$
$10^{-3}$ 0.09 23.8 2.3 2.9 257.12 $(12,1727)$
$10^{-2}$ 0.13 14.0 $\mathbf{2.0}$ $\mathbf{2.7}$ 1056.79 $(8,950)$
11 w4a $10^{-5}$ 2.36 8.5 $\mathbf{3.1}$ $\mathbf{3.1}$ 1516.61 $(5,553)$
$10^{-4}$ 2.33 10.3 $\mathbf{3.1}$ $\mathbf{3.1}$ 1655.50 $(6,701)$
$10^{-3}$ 2.15 16.2 $\mathbf{3.1}$ $\mathbf{3.1}$ 4044.59 $(8,1163)$
$10^{-2}$ 2.02 16.4 $\mathbf{3.1}$ $\mathbf{3.1}$ 2129.33 $(8,1188)$
12 w5a $10^{-5}$ 0.27 11.8 2.8 3.6 77.06 $(9,657)$
$10^{-4}$ 0.10 17.3 $\mathbf{2.1}$ 3.8 78.92 $(11,1136)$
$10^{-3}$ 0.14 19.7 $\mathbf{2.1}$ $\mathbf{3.3}$ 87.27 $(12,1059)$
$10^{-2}$ 0.10 20.5 2.5 $\mathbf{3.3}$ 77.20 $(13,1354)$

### v3

@@ 7 Conclusion
In this paper, we develop an efficient algorithm to solve the (MPEC) derived from the bilevel optimization model for hyperparameter selection for SVC. Specifically, we smoothed the complementarity constraints in the (MPEC) and proposed the smoothing damped Newton method to solve it. Under certain assumption, the proposed SDNM converges to C-stationary point under MPEC-LICQ with subproblem enjoys a quadratic convergence rate under proper assumptions. Finally, extensive numerical results on the datasets from the LIBSVM library verified the efficiency of SDNM over almost all the datasets used in this paper comparing with the other two methods. SDNM demonstrates high efficiency in practice with superior generalization performance on classification prediction and solving the over-fitting problem, and is always faster than the existing approaches.